OpenStudy (anonymous):

Can some one help me solve this? x=39 cos(23)

OpenStudy (debbieg):

I don't follow, what is there to solve? You have an expression for x. You would have to enter it in your calculator to get a numerical value for x.

OpenStudy (anonymous):

Well would I multiply divide add subtract ?

OpenStudy (debbieg):

you have 39 cos(23) That is \(\Large 39 \times \cos(23) \) so you need to multiply 39 and cos(23). I'm assuming that you probably mean 23 DEGREES, so your calculator needs to be in degree mode, not radian mode (unless you mean 23 radians).

OpenStudy (anonymous):

hold on this is the question. The distance between the end of a lamppost's shadow and the top of the lamppost is 39 feet. The angle measure of the shadow is 23°. How long is the shadow of the lamppost, rounded to the nearest tenth of a foot? Type the numeric answer only in the box below. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (debbieg):

OK, and since cos(a)=adj/hyp, you set up: cos(23)=x/39 And then you solved that for x and got : x=39*cos(23) right? That's perfect so far.

OpenStudy (anonymous):

Yeah.

OpenStudy (debbieg):

You did all the hard part. :) all that's left is to punch that up in your calculator. evaluate: 39 * cos(23)

OpenStudy (debbieg):

and round to nearest tenth (one decimal place)

OpenStudy (anonymous):

So would it be 39/23 ? 1.6956521 round and it would be 2.10 ?

OpenStudy (debbieg):

noooo..... why are you changing what you have: 39 * cos(23) to 39/23 ??? Those are completely different. You understand that cosine is a FUNCTION, right? you should have a button on your calculator that says "COS" or "COS( )", something like that. You have to EVALUATE the COSINE of the angle, 23 degrees.

OpenStudy (debbieg):

it should be a button on your calculator (if you have the kind of calculator that you need). Just like you have a square root button to evaluate a square root, you need a cosine button to evaluate cosines.

OpenStudy (anonymous):

I found one I didn't realize that sin cos tan was on it.

OpenStudy (debbieg):

If I asked you to evaluate \(\Large 39 \sqrt{23}\), you would just multiply the two, right? you wouldn't CHANGE that to \(\Large \dfrac{39}{23}\), right? :) Same idea here. \(\Large 39 \cos{(23)}\) means to multiply 39 times the cosine of the angle 23. Just punch it up. :)

OpenStudy (anonymous):

-21.1? o.O I'm not good at this at all.

OpenStudy (debbieg):

Well, it certainly shouldn't be negative... hang on....

OpenStudy (debbieg):

Yes, just what I suspected. You entered it correctly, but your calculator is set to RADIAN mode. So it is evaluating the cosine of 23 RADIANS not 23 DEGREES. :) it's important that you have the calculator in the correct mode - degrees or radians - because the answer will NOT be the same.

OpenStudy (debbieg):

Get in degree mode and re-enter it, you should have the correct answer. :)

OpenStudy (phi):

I would learn to use your calculator. In the mean time, you can type into Google 39*cos(23 deg)= to find the answer.

OpenStudy (anonymous):

How do I set it to degrees o.O

OpenStudy (debbieg):

Yeah, I considered directing her to an online calculator, but as you say, better that she understand how to use the one she has. :)

OpenStudy (debbieg):

Well, hard to know without it in front of me. If it happens to be a TI-83 or similar, look for the "mode" button. It should be under that menu. If it's anything else, get our (or google) your owner's manual.

OpenStudy (anonymous):

Oh haha found it

OpenStudy (anonymous):

so 15.23?

OpenStudy (debbieg):

hmmm. No, that isn't what I got.

OpenStudy (anonymous):

Ohh typed it in wrong 36.00?

OpenStudy (anonymous):

35.89968928

OpenStudy (debbieg):

Now I think you have it, but round to nearest TENTH... one decimal place.

OpenStudy (debbieg):

Right - that's it!

OpenStudy (anonymous):

so 36.00??

OpenStudy (anonymous):

or 36.0 because the 8 is in the 10th place right?

OpenStudy (debbieg):

I'm sorry, I had to step away and get my family lunch. :) No, round to the nearest TENTH, so look to the place one BEYOND the tenth place: \(\Large 35.8\color{red}9\) since the 100th place is a 9, you round the 10th place UP: 35.9

OpenStudy (anonymous):

Thank you so much!!